# Style Outweighs Substance: Failure Modes of LLM Judges in Alignment Benchmarking

### v2

@@ Abstract
The release of ChatGPT in November 2022 sparked an explosion of interest in post-training and an avalanche of new preference optimization (PO) methods. These methods claim superior alignment by virtue of better correspondence with human pairwise preferences, often measured by LLM-judges. In this work, we attempt to answer the following question – do LLM-judge preferences translate to progress on other, more concrete metrics for alignment, and if not, why not? We define a concrete metric for alignment, and introduce SOS-Bench (Substance Outweighs Style Benchmark), the largest standardized, reproducible LLM meta-benchmark to date. We find that (1) LLM-judge preferences do not correlate with concrete measures of safety, world knowledge, and instruction following; (2) LLM-judges have powerful implicit biases, prioritizing style over factuality and safety; and (3) the supervised fine-tuning (SFT) stage of post-training, and not the PO stage, has the greatest impact on alignment, with data scaling and prompt diversity as the driving factors. Our codebase and complete results can be found at https://github.com/penfever/sos-bench/.
@@ 3 LLM-Judge Preference Benchmarks
While the first three concerns might be expected to ameliorate over time, as LLMs become less expensive to operate, the fourth concern is foundational – there is no way for a judge to complete a preference-selection task without some degree of inductive bias (Ethayarajh et al., 2024). We observe that such a bias may be explicit (i.e., it may be introduced via the instructions to the judge) or implicit (i.e., representing a fixed value system the judge brings to the task either independently of, or in violation of, the instructions). A reasonable desideratum for an objective LLM-judge benchmark would be to make as many biases as possible explicit, and to curb the use of implicit bias. [^3] We define inductive bias as a valuable factor in machine learning, which allows the LLM to make predictions on new data based on what it learned. Implicit inductive bias we define, with some subjectivity, as any decision criteria which do not intuitively derive from the instructions provided to the judge. We note that this term is sometimes used in the ML literature to refer to the tendency of optimizers to more frequently visit some minima than others; this is not the sense in which we use the term.
In service of this goal, we devote our next section to developing an understanding of implicit bias in LLM judges.
@@ 4 Implicit bias in LLM-Judges
Experimental setting. We conduct our experiments on a series of post-trained LLAMA-3-8B base models, LLAMA-3 base without post-training, opt-125m, and several GPT checkpoints (Dubey et al., 2024; Brown et al., 2020; Zhang et al., 2022; Xu et al., 2024b; Meng et al., 2024). As of this writing, all of the checkpoints are available on HuggingFace; we provide names and references for all the post-training methods we consider in Appendix C. Our LLM-judge benchmark is Arena-Hard-Auto, from  Li et al. (2024c). We choose to make a case study of this LLM-judge benchmark because it is very popular in the literature and it makes some of the strongest claims to alignment with human preference. Unless otherwise noted, we use the standard settings for Arena-Hard-Auto, which as of this writing uses GPT-4-0314 as a baseline model and GPT-4-1106-preview as a judge. For reasons of cost, in Table 3 and Table 2, we substitute gpt-4o-mini-2024-07-18 for the standard judge. In order to conduct our experiment, we also modify the judge template. The judge template we use can be found in Section F.2. Following the authors, we report scores in the form of win-rates over a baseline model, and report pairwise preferences in the form of Likert scores.
@@ 4.1 Given explicit judgment criteria, LLM-Judges implicitly reweight them
Table 2: Judges implicitly favor stylistic criteria. When asked to render an overall judgment using a set of explicit criteria, models will implicitly prefer some of those criteria to others. We report the LLM’s overall judgment as Arena-Hard Score, alongside independent LLM judgments of five key factors in the response. Style and completeness are the strongest predictors of overall score; correctness and safety are considerably weaker, and conciseness is contra-indicated.
Method Arena-Hard Score Completeness Correctness Safety Conciseness Style
Magpie+DPO 67.76 73.88 54.43 50.95 17.02 72.79
GPT-4-0314 50.00 50.00 50.00 50.00 50.00 50.00
Meta-Llama-3-8B-Instruct 43.98 48.45 45.27 49.65 27.56 48.59
Magpie 40.81 46.65 44.21 49.60 28.19 47.15
Tulu2+DPO 38.36 44.00 43.74 49.40 28.39 44.55
GPT-4-0613 37.68 35.53 44.30 49.15 64.03 39.05
GPT-3.5-0125 34.36 30.77 44.46 49.60 65.50 36.87
WildChat 21.94 22.64 36.70 48.35 54.40 26.68
Tulu2 14.15 14.02 33.83 48.44 58.22 17.94
Bagel 08.41 09.21 28.47 47.47 56.25 11.95
Meta-Llama-3-8B 00.44 00.67 02.66 25.30 08.80 00.80
OPT-125m 00.12 00.25 02.52 22.41 07.00 00.39
In Table 2, we show the results. The scores for completeness, style, and the overall score are nearly identical; the rank order of correctness is the same as that for completeness and style, but the scores are compressed, perhaps indicating more ambivalence. Safety scores for most models are nearly identical, and conciseness is moderately anti-correlated. Given a set of criteria, LLM-judges implicitly favor completeness and style. When seen in this light, the unintuitive ranking of an 8B fine-tune above GPT-4 makes more sense; the fine-tune has produced verbose, didactic, blandly polite responses to every prompt, a set of behaviors we could collectively term stylistic reward hacking.
@@ 4.2 LLM-Judges implicitly weight criteria violations differently
In these experiments, we introduce systematic criteria violations into the model responses for the top performing model (Magpie+DPO) and recompute the model’s LLM-judge score, while leaving the rest of the pipeline unaffected. We hope to gain some indication of whether the model has understood the explicit instructions for each criteria, and how it will weight violations of those criteria. We provide samples of the altered responses in Appendix G.
@@ 5 SOS-Bench
Measuring alignment. We subdivide our benchmark into three factors (world knowledge, instruction following, and safety) and report the results for each. For the sake of comparison, we also report results from Arena-Hard-Auto. For results on individual benchmarks, we refer the reader to https://github.com/penfever/sos-bench/.
@@
Model A is more honest than model B IFF it exhibits statistically superior performance on measures of world knowledge. Although there is more to honesty than world knowledge, it is not possible for a model to intentionally tell the truth if it does not know what the truth is.
Model A is more helpful than model B IFF it exhibits statistically superior performance on measures of instruction following, because a model that correctly understands instructions is always at least as helpful as a model which fails to understand them, all else equal.
Model A is more harmless than model B IFF it exhibits statistically superior performance on measures of safety, such as red-teaming or refusal benchmarks.
@@ 6 Results
In this section, we report some interesting experimental trends which emerge on SOS-Bench.
@@
Interestingly, although all measures of alignment are positively affected by data scaling, the magnitude is much greater for instruction following (helpfulness) than for any other measure. This makes intuitive sense, given that enhancing LLM responses to declarative commands is one of the primary goals in alignment.
@@
Specialist post-training methods underperform generalist methods. We ablate the importance of prompt diversity in Table 4 by running our benchmark suite on a pool of specialist datasets designed to improve performance on one narrow task. We find that, when starting from a base checkpoint, data scaling without prompt diversity leads to overall poorer model performance compared to generalist scaling, including on specialized benchmarks.
Preference optimization trades world knowledge for improved safety and instruction following. We also conduct experiments on the second stage of post training. Somewhat surprisingly, the most significant effect we detect is a degradation in world knowledge; see Table 5. There are improvements in instruction following and safety, but they are of much smaller magnitude than the improvements during SFT. The finding that world knowledge can degrade is consistent with prior work from Bai et al. (2022), although they claim that the degradation is limited to small models. We leave the ablation of model size to future work, but note that new methods for preference optimization are often demonstrated only on small models, making this an important research consideration. There are at least two other confounds which could affect this result; the first is the choice of 2-stage dataset (we used UltraFeedback from Ding et al. (2023) for all experiments), and the choice of DPO as preference optimization algorithm. We leave the ablation of the stage-two dataset to future work, but expect the effect of this choice to be significant if it involves data scaling. We ablate the latter, and find that some other methods, most notably ORPO from Hong et al. (2024), perform better than DPO in this particular experiment; however, no method is Pareto-dominant over the baseline. See Appendix E for extended results.
All in all, we conclude that methods research in preference optimization shows marked promise, but will require more robust and careful benchmarking to instill confidence in the results.
@@ 7 Recommendations
Based on our empirical analysis, we conclude with suggestions for the research community.
Prominent researchers have urged the alignment community to converge on more precise definitions of the problem they aim to solve, as well as better measures of progress (Carlini, 2024). While there is some value in the ability to closely approximate the pairwise, undifferentiated preferences of an ’average’ human, the outsized influence of subjective design decisions such as judge instructions, inherent hackability, high unit cost, small test sets, and non-reproducibility of LLM-judges render them unsuitable for their current primary role in measuring progress in model alignment and post-training. The model alignment process necessitates trade-offs, and we cannot make those trade-offs consciously when undocumented implicit biases dominate our key metrics. We call on the research community to develop more targeted benchmarks for specific HHH factors of interest, such as the recent IFEval (Zhou et al., 2023b) for instruction following and FLASK (Ye et al., 2024) for particular skill sets, and recommend that reviewers insist that authors who wish to make broad claims about their method support those claims with general-purpose alignment benchmarks such as SOS-Bench.
We also call on the research community to move beyond the Bradley-Terry model, towards more sophisticated approaches to post-training. The fact that one model’s output is generally preferred to another model’s output does not necessarily indicate that its output is more aligned in general, and general alignment is at best an intermediate goal in any case. Naive assumptions of universal preference fail to capture the diversity between, and idiosyncrasy within, groups. Narrow LLM-judge benchmarks, with carefully curated questions, designed to evaluate specific desirable factors in isolation on particular knowledge domains, would be a useful step in this direction.
@@ 8 Related Work
A series of recent works have offered various critiques on the use of LLM-judges (Chen et al., 2024; Bavaresco et al., 2024; Wei et al., 2024). Several known confounds for pairwise benchmarks are well-established in the literature, in particular a bias by both humans and LLMs in favor of longer responses and a preference of LLMs for the style of their own output; efforts have been made to control for length as a confound (Panickssery et al., 2024; Park et al., 2024; Dubois et al., 2024; Wu & Aji, 2023). There also exists considerable prior literature critiquing human judgment bias and proposing mechanisms for collective decision making, including social choice theory and prospect theory (Ge et al., 2024; Ethayarajh et al., 2024). Finally, there exists a literature on the limitations of RLHF and pairwise preferences, including findings that universal AI alignment using RLHF is impossible (Singhal et al., 2024; Casper et al., 2023; Lambert & Calandra, 2024; Mishra, 2023; Lambert et al., 2023). Extending prior work, our research focuses on novel semantic biases, and also introduces the concept of an implicit weighting of factors on the part of LLM judges. Finally, our work draws on all of the aforementioned traditions to produce a meta-analysis on the progress, or lack thereof, of methods for LLM post-training.
@@
Similar to some aspects of our work is the excellent contribution of Ye et al. (2024), who propose judging a fine-grained evaluation of LLM skills instead of coarse preferences. Unlike their work, our work reports both fine and coarse preferences, and uses the explicit fine-grained preferences to reveal the implicit LLM bias in determining coarse preference.
@@ Appendix B Compute and Resources
The experiments for this work were conducted using NYU’s Greene cluster and on Google Cloud, with additional financial support provided by Arthur.AI.
@@
We primarily utilize the OpenAI API, following Li et al. (2024c); The cost to produce Table 2 was \$40 USD, as we replaced the GPT-4 judge with a GPT-4o-mini judge. The cost to produce Table 3 was \$25 USD, for the same reason. The cost to produce Figure 2 was \$750 USD, as we used the original GPT-4 judge so as to remain consistent with Li et al. (2024c). The cost of Table 5 was \$90.

### v3

@@ Abstract
The release of ChatGPT in November 2022 sparked an explosion of interest in post-training and an avalanche of new preference optimization (PO) methods. These methods claim superior alignment by virtue of better correspondence with human pairwise preferences, often measured by LLM-judges. In this work, we attempt to answer the following question – do LLM-judge preferences translate to progress on other, more concrete metrics for alignment, and if not, why not? We define a concrete metric for alignment, and introduce SOS-Bench (Substance Outweighs Style Benchmark), the largest standardized, reproducible LLM meta-benchmark to date. We find that (1) LLM-judge preferences do not correlate with concrete measures of safety, world knowledge, and instruction following; (2) LLM-judges have powerful implicit biases, prioritizing style over factuality and safety; and (3) the supervised fine-tuning (SFT) stage of post-training has a large impact on alignment, with data scaling and prompt diversity as the driving factors. Our codebase and complete results can be found at https://github.com/penfever/sos-bench.
@@ 3 LLM-Judge Preference Benchmarks
The Arena-Hard-Auto Pipeline. For convenience, we summarize the steps in the arena-hard-auto judgment pipeline here. The chosen LLM judge conducts a pairwise comparison against a baseline model (GPT-4-0314 in the original paper), scoring outputs on 5-point Likert scale. The judge generates its own response to the question before judging, and uses chain-of-thought prompting for consistent judgments. The paper implements a two-game setup to avoid position bias, aggregating 1000 judgments per model using Bradley-Terry, resulting in final scores and confidence intervals through bootstrapping. This judgment pipeline has been shown to have strong correlation with the judgments of humans Li et al. (2024c) 95% confidence intervals are reported as part of the benchmark; these can be as high as 4% for judgments close to the 50th percentile score. However, we believe that these uncertainty estimates are too conservative – therefore, in Section I.3, we independently ablate these choices for a single judge, and find that the choice of baseline and judge template are the most impactful factors. Surprisingly, we find that replacing the carefully filtered Arena-hard questions with questions from popular non-technical subreddits such as AskHistorians has little effect on the ranking, suggesting that LLM judge scores are produced in a manner only modestly dependent on the knowledge domain.
@@
While the first three concerns might be expected to ameliorate over time, as LLMs become less expensive to operate, the fourth concern is foundational – there is no way for a judge to complete a preference-selection task without some degree of inductive bias (Ethayarajh et al., 2024). We observe that such a bias may be explicit (i.e., it may be introduced via the instructions to the judge) or implicit (i.e., representing a fixed value system the judge brings to the task either independently of, or in violation of, the instructions). A reasonable desideratum for an objective LLM-judge benchmark would be to make as many biases as possible explicit, and to curb the use of implicit bias. [^3] We define inductive bias as a valuable factor in machine learning, which allows the LLM to make predictions on new data based on what it learned. Implicit inductive bias we define, with some subjectivity, as any decision criteria which do not intuitively derive from the instructions provided to the judge. We note that this term is sometimes used in the ML literature to refer to the tendency of optimizers to more frequently visit some minima than others; this is not the sense in which we use the term. In service of this goal, we devote our next section to developing an understanding of implicit bias in LLM judges.
@@ 4 Implicit bias in LLM-Judges
Experimental setting. We conduct our experiments on a series of post-trained LLAMA-3-8B base models, LLAMA-3 base without post-training, opt-125m, and several GPT checkpoints (Dubey et al., 2024; Brown et al., 2020; Zhang et al., 2022; Xu et al., 2024b; Meng et al., 2024). As of this writing, all of the checkpoints are available on HuggingFace; we provide names and references for all the post-training methods we consider in Appendix C. Our LLM-judge benchmark is Arena-Hard-Auto, from  Li et al. (2024c). We choose to make a case study of this LLM-judge benchmark because it is very popular in the literature and it makes some of the strongest claims to alignment with human preference. Unless otherwise noted, we use the standard settings for Arena-Hard-Auto, which as of this writing uses GPT-4-0314 as a baseline model and GPT-4-1106-preview as a judge. For reasons of cost, in Table 3, we substitute gpt-4o-mini-2024-07-18 for the standard judge. In order to conduct our experiment, we also modify the judge template. The judge template we use can be found in Section F.2. Following the authors, we report scores in the form of win-rates over a baseline model, and report pairwise preferences in the form of Likert scores.
@@ 4.1 Given explicit judgment criteria, LLM-Judges implicitly reweight them
Figure 2: Judges implicitly reweight explicit criteria. When asked to render an overall judgment using a set of explicit criteria, models will implicitly weight some of those criteria more than others. We report the LLM’s overall judgment as Arena-Hard Score, alongside independent LLM judgments of five key factors in the response. Style is perfectly correlated with the overall score (Pearson’s R).
Table 2: The reweighting is stable across judges. We query a panel of four judges (gpt-3.5-turbo-1106, gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, claude-3-5-sonnet-20241022) and find that style predicts overall score almost perfectly for every judge in our panel (we report the average correlation and standard deviation).
Factor Spearman Pearson Std
Style 1 0.999 0
Completeness 0.964 0.963 0.04
Correctness 0.906 0.881 0.079
Safety 0.827 0.727 0.147
Conciseness 0.097 0.114 0.128
In Figure 2, we show that the style score correlates perfectly with the overall score. By contrast, safety scores for most models are nearly identical, and conciseness is only weakly correlated. Given a set of criteria, LLM-judges implicitly favor completeness and style, and this behavior is strongly conserved across judges: see Table 2. When seen in this light, the unintuitive ranking of an 8B fine-tune above GPT-4 makes more sense; the fine-tune has produced verbose, didactic, blandly polite responses to every prompt, a set of behaviors we could collectively term stylistic reward hacking. It is important to note that factors are not necessarily independent; we analyze the degree and direction of cross-correlation using factor analysis in Appendix K.
@@ 4.2 LLM-Judges implicitly weight criteria violations differently
In these experiments, we introduce systematic criteria violations into the model responses for the top performing model (Magpie+DPO) and recompute the model’s LLM-judge score, while leaving the rest of the pipeline unaffected. We then report the loss, expressed as the percentage difference between the base score and the new score. We hope to gain some indication of whether the model has understood the explicit instructions for each criteria, and how it will weight violations of those criteria. We provide samples of the altered responses in Appendix G.
@@ 5 SOS-Bench
Measuring alignment. We subdivide our benchmark into three factors (world knowledge, instruction following, and safety) and report the results for each. For the sake of comparison, we also report results from Arena-Hard-Auto. For results on individual benchmarks, we refer the reader to https://github.com/penfever/sos-bench.
@@
Model A is more honest than model B IFF it exhibits statistically superior performance on measures of world knowledge. Although there is more to honesty than world knowledge, it is not possible for a model to tell the truth if it does not know what the truth is. Model A is more helpful than model B IFF it exhibits statistically superior performance on measures of instruction following, because a model that correctly understands instructions is always at least as helpful as a model which fails to understand them, all else equal. Model A is more harmless than model B IFF it exhibits statistically superior performance on measures of safety, such as red-teaming or refusal benchmarks.
@@ 6 Results
Interestingly, although all measures of alignment are positively affected by data scaling, the magnitude is much greater for instruction following (helpfulness). This makes intuitive sense, given that enhancing LLM responses to declarative commands is one of the primary goals in alignment.
@@
Specialist post-training methods underperform generalist methods. We ablate the importance of prompt diversity in Table 4 by running our benchmark suite on a pool of specialist datasets designed to improve performance on one narrow task. We find that, when starting from a base checkpoint, data scaling without prompt diversity leads to overall poorer model performance compared to generalist scaling, including on specialized benchmarks. We define prompt diversity as the breadth of domain knowledge and the variation in semantic structures among prompts.
Preference optimization trades world knowledge for improved safety and instruction following. We also conduct experiments on the second stage of post training. Somewhat surprisingly, the most significant effect we detect is a degradation in world knowledge; see Table 5. There are improvements in instruction following and safety, but they are of much smaller magnitude than the improvements during SFT. The finding that world knowledge can degrade is consistent with prior work from Bai et al. (2022), although they claim that the degradation is limited to small models. We leave the ablation of model size to future work, but note that new methods for preference optimization are often demonstrated only on small models, making this an important research consideration. There are at least two other confounds which could affect this result; the first is the choice of 2-stage dataset (we used UltraFeedback from Ding et al. (2023) for all experiments), and the choice of DPO as preference optimization algorithm. We leave the ablation of the stage-two dataset to future work, but expect the effect of this choice to be significant if it involves data scaling. We ablate the latter, and find that some other methods, most notably ORPO from Hong et al. (2024), perform better than DPO in this particular experiment; however, no method is Pareto-dominant over the baseline. See Appendix E for extended results. All in all, we conclude that methods research in preference optimization shows marked promise, but will require more robust and careful benchmarking to instill confidence in the results.
@@ 7 Recommendations
Prominent researchers have urged the alignment community to converge on more precise definitions of the problem they aim to solve, as well as better measures of progress (Carlini, 2024). While there is some value in the ability to closely approximate the pairwise, undifferentiated preferences of an ’average’ human, the outsized influence of subjective design decisions such as judge instructions, inherent hackability, high unit cost, small test sets, and non-reproducibility of LLM-judges render them unsuitable for their current primary role in measuring progress in model alignment and post-training. The model alignment process necessitates trade-offs, and we cannot make those trade-offs consciously when undocumented implicit biases dominate our key metrics. We call on the research community to develop more targeted benchmarks for specific HHH factors of interest, such as the recent IFEval (Zhou et al., 2023b) for instruction following and FLASK (Ye et al., 2024) for particular skill sets, and recommend that reviewers insist that authors who wish to make broad claims about their method support those claims with general-purpose alignment benchmarks such as SOS-Bench. We also call on the research community to move beyond the Bradley-Terry model, towards more sophisticated approaches to post-training.
@@ 8 Related Work
A series of recent works have offered various critiques on the use of LLM-judges (Chen et al., 2024; Bavaresco et al., 2024; Wei et al., 2024; Sorensen et al., 2024). Several known confounds for pairwise benchmarks are well-established in the literature, in particular a bias by both humans and LLMs in favor of longer responses and a preference of LLMs for the style of their own output; efforts have been made to control for length as a confound (Panickssery et al., 2024; Park et al., 2024; Dubois et al., 2024; Wu & Aji, 2023). There also exists considerable prior literature critiquing human judgment bias and proposing mechanisms for collective decision making, including social choice theory and prospect theory (Ge et al., 2024; Ethayarajh et al., 2024). Finally, there exists a literature on the limitations of RLHF and pairwise preferences, including findings that universal AI alignment using RLHF is impossible (Singhal et al., 2024; Casper et al., 2023; Lambert & Calandra, 2024; Mishra, 2023; Lambert et al., 2023). Extending prior work, our research focuses on novel semantic biases, and also introduces the concept of an implicit weighting of factors on the part of LLM judges. Finally, our work draws on all of the aforementioned traditions to produce a meta-analysis on the progress, or lack thereof, of methods for LLM post-training.
@@
Similar to some aspects of our work is the excellent contribution of Ye et al. (2024), who propose judging a fine-grained evaluation of LLM skills instead of coarse preferences. We report both fine and coarse preferences, and use the explicit fine-grained preferences to reveal the implicit LLM bias in determining coarse preference.
@@ 9 Impact / Limitations
10 Reproducibility Statement
We have tried to ensure that the research described in this paper is reproducible. Our benchmark comparisons in Section 3 and Table 1 can be reproduced using the LiveBench, ChatBot Arena, Arena-Hard-Auto, HELM and BenchBench leaderboards, which are publicly available. The checkpoints used in our experiments in Section 4 will be made available once our work is de-anonymized; our chat templates and experimental details are in the appendix. Our repository and code for Section 5 is already publicly available, and our list of benchmarks is documented in Table 8 as well.
@@ Appendix B Compute and Resources
We primarily utilize the OpenAI API, following Li et al. (2024c); The cost to produce Figure 2 was \$40 USD, as we replaced the GPT-4 judge with a GPT-4o-mini judge. The cost of Table 2 was \$600 USD. The cost to produce Table 3 was \$25 USD, for the same reason. The cost to produce Figure 3 was \$750 USD, as we used the original GPT-4 judge so as to remain consistent with Li et al. (2024c). The cost of Table 5 was \$90.
@@ Appendix C Method Comparison
Code Llama 3 from ajibawa-2023 (2024) 2024 1 N/A N/A
Numina Math CoT from LI et al. (2024) 2024 1 N/A N/A
Tulu Human from Ivison et al. (2023) 2024 1 N/A N/A
@@ I.2 Hyperparameter Ablations
I.3 Arena-Hard-Auto Ablations
Using GPT-4o-mini as a judge, we progressively ablate the design choices in arena-hard-auto.
We find that several design choices, such as the template instruction for the judge to answer the question itself before responding (noselfref) and judging all pairings twice with independent ordering (nopairwise) have very little effect on the judgments, despite dramatically increasing the cost of a single run.
Considerably more impactful are the particular instructions given to the model (template) and the choice of baseline model to compare against.
Surprisingly, we find that changing the questions (questions) does not have a particularly strong effect on the rank order of models, suggesting that LLM judge scores are produced in a manner largely independent of the question domain. This is potentially problematic for domains which require either highly specialized knowledge, such as science and medicine, or require distinct stylistic conventions to be followed, such as law and creative writing.
Figure 5: Comparative impact of changes to the Arena-Hard-Auto methodology. The last row represents the strength of correlation between the ablation and the base (Pearson’s R). We observe the highest sensitivity when changing the baseline model used for pairwise comparisons. Interestingly, changing the questions does not have a particularly strong effect on the rank order of models, suggesting that what LLM judges measure is not particularly attuned to subject matter.